Solve. Identify all double roots.
8y^2 + 1 = 6y

Answer by mananth(16946)

8y^2 + 1 = 6y
8y^2 - 6y+1=0
8y^2-4y-2y+1=0
4y(2y-1)-1(2y-1)=0
(2y-1)(4y-1)=0
y=1/2 OR y=1/4